OpenStudy (anonymous):

Rust, chemically, is iron(III) oxide,Fe2O3.On a large ship, 16kg of rust accumulates in a day. Calculate how many kilograms of iron and how many kilograms of oxygen would be used up. ( Ans:Fe=4.19 O=.9024 . Give me details)

OpenStudy (michele_laino):

hint: the molar mass of \(Fe_2O_3\) is \(159.70\) So, the number of mol of \(Fe_2O_3\) is: \[n = \frac{{16000}}{{159.70}} = ...?\]

OpenStudy (anonymous):

n=160/159.70=1

OpenStudy (michele_laino):

hint: 16 Kg are equivalent to 16,000 grams, so we have to do this computation: \[n = \frac{{16000}}{{159.70}} = ...?\]

OpenStudy (anonymous):

n=100.18

OpenStudy (michele_laino):

correct!

OpenStudy (michele_laino):

now, 1 mol of \(Fe_2O_3\) contains 2 moles of \(F_e\) and 3 moles of \(O\), so we can write: \[\Large \begin{gathered} {m_{Fe}} = \frac{{100.18 \times 2 \times 55.85}}{{1000}} = ...{\text{Kg}} \hfill \\ \hfill \\ {m_O} = \frac{{100.18 \times 3 \times 16}}{{1000}} = ...{\text{Kg}} \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

where \(m_O\) and \(m_{Fe}\) are the masses of oxygen and iron respectively

OpenStudy (anonymous):

mFe=11.198Kg & mO=4.8

OpenStudy (anonymous):

Thanks a lot for solving this

OpenStudy (michele_laino):

that's right! As we can see if we compute this sum: \(11.2+4.8\), we get the initial mass, namely \(16\) Kg

OpenStudy (michele_laino):

so, the mass is conserved, according to the law of \(Lavoisier\)